Last week Anna ran 20 miles. This week she ran 25 miles. The increase is what percent of the number of miles she ran last week?
Answer by fractalier(2101)

Percentage increase = [(New - Old) / Old] x 100%
so if we plug in we get
% increase = [(25 - 20) / 20] x 100% = (5/20) x 100% = 25%